OpenStudy (anonymous):

"With y= x^2 - 2 , no matter what value you enter, you will never get a result less than -2 ." Why is that?

OpenStudy (anonymous):

because for all values of x, both positive and negative, x^2 is greater than or equal to 0

OpenStudy (anonymous):

for example, if x = -1, x^2 = 1 so y = 1-2 = -1 which is greater than -2 if x = 0, x^2 =0 y = 0-2 = -2 which is equal to -2 if x =1, x^2 = 1 y = 1-2 = -1 which is greater than -2